# Embedding Inversion via Conditional Masked Diffusion Language Models

### v1

@@ Abstract
We frame embedding inversion as conditional masked diffusion, recovering all tokens in parallel through iterative denoising rather than sequential autoregressive generation. A masked diffusion language model is conditioned on the target embedding via adaptive layer normalization, requiring only 8 forward passes through a 78M parameter model with no access to the target encoder. On 32-token sequences across three embedding models, the method achieves 81.3% token accuracy and 0.87 cosine similarity. Source code and live demo are available at https://github.com/hanxiao/embedding-inversion-demo.
@@ 1 Introduction
We propose an alternative formulation: embedding inversion as conditional masked diffusion. Starting from a fully masked sequence, a denoising model iteratively reveals tokens at all positions in parallel, conditioned on the target embedding vector via adaptive layer normalization. The key structural difference is that correction is built into the diffusion process itself: each denoising step refines all positions simultaneously using global context, without ever reembedding the current hypothesis. This eliminates the need for access to the target encoder at inference time and reduces attack cost to a fixed number of forward passes through a small model of 78M parameters.
The approach is encoder agnostic by construction. The embedding vector enters only through AdaLN modulation of layer normalization parameters, so the same architecture and training procedure applies to any embedding model without alignment training or architecture specific modifications. We demonstrate this by training on three different encoders: jina-embeddings-v3 with 1024 dimensions, Qwen3-Embedding-0.6B with 1024 dimensions, and EmbeddingGemma-300m with 768 dimensions.
@@
We present the first application of masked diffusion language models to embedding inversion, replacing autoregressive generation and iterative reembedding with parallel denoising. The approach is encoder agnostic: we train on three embedding models without alignment training or architecture specific modifications. We systematically compare four decoding strategies, identifying adaptive remasking during Euler sampling as the best quality-efficiency trade-off for parallel generation. On 32-token sequences, our 78M parameter model recovers up to 81.3% of tokens with 0.87 cosine similarity from a single embedding vector, requiring no access to the target encoder at inference time.
@@ 3 Method
We use the following notation throughout: $\mathbf{x}=(x_{1},\ldots,x_{n})$ denotes a token sequence of length $n$ from vocabulary $\mathcal{V}$; $\mathbf{e}\in\mathbb{R}^{d}$ denotes the embedding vector; $t\in[0,1]$ denotes the diffusion timestep with $t=0$ being fully unmasked and $t=1$ being fully masked; $\theta$ denotes the model parameters; $\mathbf{c}\in\mathbb{R}^{D_{h}}$ denotes the projected conditioning vector with hidden dimension $D_{h}=768$; $x_{t}$ denotes the masked sequence at timestep $t$; $x_{0}$ denotes the original unmasked sequence.
@@ 3.2 Masked Diffusion Process
where $\mathcal{D}$ is the data distribution, the sum is over masked positions only, and the $1/t$ weighting prioritizes early timesteps with more masked tokens. This weighting emphasizes learning global structure over local refinements.
@@ 3.3 Model Architecture
Our model consists of three components: embedding projection, transformer backbone, and adaptive layer normalization conditioning (Figure 1). The input embedding $\mathbf{e}\in\mathbb{R}^{d}$ is projected to the transformer hidden dimension $D_{h}=768$ via a two-layer MLP:
@@
where $\mathbf{W}_{1}\in\mathbb{R}^{D_{h}\times d}$, $\mathbf{W}_{2}\in\mathbb{R}^{D_{h}\times D_{h}}$, and $\mathbf{b}_{1},\mathbf{b}_{2}\in\mathbb{R}^{D_{h}}$ are learned parameters. We use an 8-layer transformer with $D_{h}=768$ hidden dimensions, 12 attention heads, and FFN dimension 3072. Input and output embeddings are weight-tied to reduce parameters given the large vocabulary size $|\mathcal{V}|=50257$.
@@
where $\mathbf{h}^{(\ell)}\in\mathbb{R}^{n\times D_{h}}$ is the input to layer $\ell$, $\mu(\cdot)$ and $\sigma(\cdot)$ compute mean and standard deviation over the hidden dimension, and $\odot$ denotes element-wise multiplication. This formulation allows the conditioning signal and timestep to independently modulate the layer normalization at each depth, providing fine-grained control over feature representations. The complete model has approximately 270M parameters due to the large vocabulary embeddings, but only 78M trainable parameters consisting of the 8 transformer layers, embedding projection MLP, and AdaLN conditioning MLPs.
@@ 4 Experimental Results
We train on 2M samples from C4 (Raffel et al., 2020), filtered to 32 tokens. We use the GPT-2 tokenizer with vocabulary size 50,257. Training uses batch size 400 for 200K steps with AdamW optimizer at learning rate $10^{-4}$ and EMA decay 0.9999. We employ a log-linear noise schedule with $\lambda=5.0$ following Sahoo et al. (2024). Timesteps are sampled uniformly from $[0,1]$. Embeddings are computed using the target encoder and cached. We evaluate on three embedding models with different architectures and dimensionalities: jina-embeddings-v3 (Sturua et al., 2024) with 570M parameters and 1024-dimensional embeddings, Qwen3-Embedding-0.6B with 600M parameters and 1024-dimensional embeddings, and EmbeddingGemma-300m with 300M parameters and 768-dimensional embeddings. We train separate models for each encoder using multilingual data from mC4 to assess generalization across embedding spaces.
We compare four decoding strategies at inference time. Sequential greedy decoding iteratively unmasks tokens left to right by taking $x_{i}=\arg\max_{v\in\mathcal{V}}p_{\theta}(v|x_{<i},[\text{MASK}]^{n-i},\mathbf{e},t)$ where t=(ni)/nt=(n-i)/n corresponds to the fraction of remaining masked tokens, producing highly coherent text through left-to-right generation but sacrificing the parallel nature of diffusion. Euler sampling uses the Euler method for the reverse diffusion process, starting from $x_{1}=[\text{MASK}]^{n}$ and taking uniform timesteps from $t=1$ to $t=0$, sampling from $p_{\theta}(x_{0,i}|x_{t},t,\mathbf{e})$ for all positions simultaneously. Euler with remasking re-masks positions where $\max_{v}p_{\theta}(v|x_{t},t,\mathbf{e})<\tau$ after each Euler step, refining low-confidence predictions in subsequent steps. Two-stage decoding combines sequential and parallel approaches by first generating a hypothesis via sequential greedy decoding, then refining it using Euler sampling initialized at this hypothesis. We use token accuracy, exact match, cosine similarity, BLEU, and perplexity under GPT-2 as evaluation metrics.
4.1 Performance Across Encoders
@@
Table 2 compares four decoding strategies across all three encoders on 10 languages. Cosine similarity is averaged over the same sentence translated into English, Chinese, German, Japanese, French, Spanish, Korean, Russian, Arabic, and Portuguese. Sequential greedy consistently achieves the highest similarity across encoders. Qualitative examples of decoded text across languages and decoding strategies are provided in Appendix 69.
@@ 5 Conclusion
We presented embedding inversion via conditional masked diffusion, achieving 81.3% token accuracy across three embedding models with a 78M parameter decoder that requires no access to the target encoder. As inversion methods evolve from architecture specific to fully agnostic, embeddings should be treated as sensitive data requiring protection equivalent to the original text.
Appendix A Implementation Details
A.1 Hyperparameters
Table 4: Complete hyperparameter configuration.
Parameter Value
Transformer layers 8
Hidden dimension 768
Attention heads 12
MLP dimension 3072
Dropout 0.1
Batch size 400
Learning rate $10^{-4}$
Weight decay 0.01
EMA decay 0.9999
Noise schedule $\lambda$ 5.0
Sequence length 32
Training steps 200K
A.2 Computational Requirements
Training on a single A100 GPU takes approximately 48 hours for 200K steps. Inference using sequential decoding takes 150ms per sequence on the same hardware. Euler sampling is significantly faster at approximately 50ms per sequence due to parallel token prediction, though it achieves slightly lower quality. Memory requirements during training peak at 24GB including optimizer states and activations for the batch size of 400. The model checkpoint size is 312MB including both model parameters and EMA weights.
Table 5: Performance milestones during training for jina-v3 encoder. Sequential greedy decoding.
Training Steps Token Acc. Exact Match Cosine Sim. BLEU
10K 40.2% 1.3% 0.65 22.1
25K 58.7% 5.8% 0.76 35.4
50K 69.7% 12.3% 0.83 45.2
100K 70.8% 13.7% 0.84 46.8
200K 71.2% 14.1% 0.84 47.3
Appendix B Qualitative Examples
Tables 69 show qualitative inversion examples. Table 6 compares four decoding strategies on the same English input using jina-embeddings-v3. Tables 7, 8, and 9 show sequential greedy decoding across six languages for each encoder.
Table 6: Decoding strategy comparison on jina-embeddings-v3. Input: “The advancement of artificial intelligence has fundamentally transformed modern society.”
Method Recovered Text Sim.
Sequential The revolution of artificial society is a new technology that is the most advanced in the world. The goal of the world is to increase the lives of the people 0.82
Euler The imagination that we transformed and automitting technology has become optimized by the modern Intelligence Agency that the world has endured in revolutionary technologies that 0.80
Euler + Remask Artificial is the future of the outbreak of mischievous technology. Many people around the world are analyzed by transforming society by life, of 0.78
Two-Stage Today, the AI industry is one of the the sight revolutions, a history of Modern and Advanced Languages. He has brought the world of Artificial 0.79
Table 7: Multilingual inversion examples using jina-embeddings-v3 with sequential greedy decoding.
Lang Original Recovered Sim.
en The advancement of artificial intelligence has fundamentally transformed modern society. The revolution of artificial society is a new technology that is the most advanced in the world. 0.82
zh 人工智能的进步从根本上改变了现代社会解决复杂问题的方式。 The complexity of the modern life changes in complexity. The problem is that the problem is to create a new life in the world. 0.65
de Die Fortschritte in der künstlichen Intelligenz haben die moderne Gesellschaft grundlegend verändert. The modern world of the modern society is increasingly changing in the world of the modern life. 0.69
ja 人工知能の進歩は現代社会における問題解決の方法を根本的に変革しました。 The modern life changes in the world of the problem. The problem of the problem is to transform the complexity of the artificial intelligence. 0.76
fr Les progrès de l’intelligence artificielle ont fondamentalement transformé la société moderne. The transformation of the modern society is the most important factor in the world. 0.70
es Los avances en inteligencia artificial han transformado fundamentalmente la sociedad moderna. The modern revolution of the artificial society is a major transformation in the United States. 0.81
Table 8: Multilingual inversion examples using Qwen3-Embedding with sequential greedy decoding.
Lang Original Recovered Sim.
en The advancement of artificial intelligence has fundamentally transformed modern society. The most widely accepted and transformative of the natural and social media in the contemporary and technological aspects of the entire life. 0.78
zh 人工智能的进步从根本上改变了现代社会解决复杂问题的方式。 The world has become more complex than ever before, as a result of the real problem of the way, the new ways of the artificial intelligence. 0.65
de Die Fortschritte in der künstlichen Intelligenz haben die moderne Gesellschaft grundlegend verändert. The real-life of the German civilization is transforming the world into the technology of the German civilization. 0.65
ja 人工知能の進歩は現代社会における問題解決の方法を根本的に変革しました。 The world of human society is increasingly important in the ways of the human society as the result of the new generation. 0.57
fr Les progrès de l’intelligence artificielle ont fondamentalement transformé la société moderne. The real life of the society is the increase of the the chemical in the the et aux ité de la la de la de la. 0.65
es Los avances en inteligencia artificial han transformado fundamentalmente la sociedad moderna. The real world of the way the technology is transforming the average of the natural and social media in the entire world. 0.67
Table 9: Multilingual inversion examples using EmbeddingGemma with sequential greedy decoding.
Lang Original Recovered Sim.
en The advancement of artificial intelligence has fundamentally transformed modern society. The systems of the most important to the human intelligence, the most important to the human sector of the modern approach. 0.68
zh 人工智能的进步从根本上改变了现代社会解决复杂问题的方式。 A complex system is a complex approach to the systems of the ability to develop the most powerful and the ability to develop. 0.61
de Die Fortschritte in der künstlichen Intelligenz haben die moderne Gesellschaft grundlegend verändert. Die Forschung ist die Zeit der Zeit der Entwicklung von der Zeit der Entwicklung der Entwicklung der Zeit von der Welt. 0.57
ja 人工知能の進歩は現代社会における問題解決の方法を根本的に変革しました。 A system of the systems of the human development of the importance of the current and the ability to develop the cell-based approach. 0.59
fr Les progrès de l’intelligence artificielle ont fondamentalement transformé la société moderne. La technologie de la série de la société de la planète, de la gestion de la planète, de la science. 0.62
es Los avances en inteligencia artificial han transformado fundamentalmente la sociedad moderna. La ingeniería de la historia de la inteligencia de la organización de la atención de la que las tecnologías. 0.64

### v2

@@ Abstract
We frame embedding inversion as conditional masked diffusion, recovering all tokens in parallel through iterative denoising rather than sequential autoregressive generation. A masked diffusion language model is conditioned on the target embedding via adaptive layer normalization, requiring only 8 forward passes through a 78M parameter model with no access to the target encoder. On 32-token sequences across three embedding models, the method achieves up to 81.3% token accuracy. Source code and live demo are available at https://github.com/jina-ai/embedding-inversion-demo.
@@ 1 Introduction
We propose an alternative formulation: embedding inversion as conditional masked diffusion. Starting from a fully masked sequence, a denoising model iteratively reveals tokens at all positions in parallel, conditioned on the target embedding vector via adaptive layer normalization. Correction is built into the diffusion process itself: each step refines all positions simultaneously using global context, without ever reembedding the current hypothesis. The embedding vector enters only through AdaLN modulation, making the approach encoder agnostic: the same architecture applies to any embedding model without alignment training. We validate on three encoders with different architectures and dimensionalities, achieving up to 81.3% token recovery on 32-token sequences with no access to the target encoder at inference time.
@@ 3 Method
We use the following notation throughout: $\mathbf{x}=(x_{1},\ldots,x_{n})$ denotes a token sequence of length $n$ from vocabulary $\mathcal{V}$; $\mathbf{e}\in\mathbb{R}^{d}$ denotes the embedding vector; $t\in[0,1]$ denotes the diffusion timestep with $t=0$ being fully unmasked and $t=1$ being fully masked; $\theta$ denotes the model parameters; $\mathbf{c}\in\mathbb{R}^{D_{h}}$ denotes the projected conditioning vector with hidden dimension $D_{h}$; $x_{t}$ denotes the masked sequence at timestep $t$; $x_{0}$ denotes the original unmasked sequence.
@@ 3.2 Masked Diffusion Process
where $\mathcal{D}$ is the data distribution, the sum is over masked positions only, and the $1/t$ weighting upweights the low-noise regime ($t\to 0$), where few tokens remain masked and precise reconstruction matters most.
@@ 3.3 Model Architecture
Our model consists of three components: embedding projection, transformer backbone, and adaptive layer normalization conditioning (Figure 1). The input embedding $\mathbf{e}\in\mathbb{R}^{d}$, where $d$ is determined by the target encoder, is projected to the transformer hidden dimension $D_{h}$ via a two-layer MLP:
@@
where $\mathbf{W}_{1}\in\mathbb{R}^{D_{h}\times d}$, $\mathbf{W}_{2}\in\mathbb{R}^{D_{h}\times D_{h}}$, and $\mathbf{b}_{1},\mathbf{b}_{2}\in\mathbb{R}^{D_{h}}$ are learned parameters. We use an 8-layer transformer with 12 attention heads and FFN dimension $4D_{h}$. In our experiments, $D_{h}=768$, yielding FFN dimension 3072. Input and output embeddings are weight-tied to reduce parameters given the large vocabulary size $|\mathcal{V}|=50257$.
@@
where $\mathbf{h}^{(\ell)}\in\mathbb{R}^{n\times D_{h}}$ is the input to layer $\ell$, $\mu(\cdot)$ and $\sigma(\cdot)$ compute mean and standard deviation over the hidden dimension, and $\odot$ denotes element-wise multiplication. This formulation allows the conditioning signal and timestep to independently modulate the layer normalization at each depth, providing fine-grained control over feature representations.
3.4 Decoding Strategies
We consider four strategies for generating tokens from the trained model.
Sequential greedy decoding unmasks tokens left to right:
$x_{i}=\arg\max_{v\in\mathcal{V}}p_{\theta}(v\mid x_{<i},[\text{MASK}]^{n-i},\mathbf{e},t)$ (11)
where t=(ni)/nt=(n-i)/n is the fraction of remaining masked tokens. This produces coherent text but sacrifices the parallel nature of diffusion.
Euler sampling applies the Euler method to the reverse diffusion process, starting from $x_{1}=[\text{MASK}]^{n}$ with uniform timesteps from t=1t{=}1 to t=0t{=}0:
$\hat{x}_{0,i}\sim p_{\theta}(x_{0,i}\mid x_{t},t,\mathbf{e})\quad\forall\,i$ (12)
sampling all positions simultaneously at each step.
Euler with remasking adds a correction mechanism: after each Euler step, a fraction $\tau$ of positions with the lowest confidence $\max_{v}p_{\theta}(v\mid x_{t},t,\mathbf{e})$ are re-masked:
[…]
allowing subsequent steps to refine uncertain predictions. We find $\tau=0.05$ optimal (Table 3).
Two-stage decoding first generates a hypothesis via sequential greedy decoding, then refines it using Euler sampling initialized at this hypothesis rather than a fully masked sequence.
@@ 4 Experimental Results
We train on 2M samples from C4 (Raffel et al., 2020), filtered to 32 tokens. We use the GPT-2 tokenizer with vocabulary size 50,257. Training uses batch size 400 for 200K steps with AdamW optimizer at learning rate $10^{-4}$ and EMA decay 0.9999. We employ a log-linear noise schedule with $\lambda=5.0$ following Sahoo et al. (2024). Timesteps are sampled uniformly from $[0,1]$. Embeddings are computed using the target encoder and cached. We evaluate on three embedding models with different architectures and dimensionalities: jina-embeddings-v3 (Sturua et al., 2025) with 570M parameters and 1024-dimensional embeddings, Qwen3-Embedding-0.6B with 600M parameters and 1024-dimensional embeddings, and EmbeddingGemma-300m with 300M parameters and 768-dimensional embeddings. We train separate models for each encoder using multilingual data from mC4 to assess generalization across embedding spaces.
The complete model has approximately 270M parameters due to the large vocabulary embeddings, but only 78M trainable parameters consisting of the 8 transformer layers, embedding projection MLP, and AdaLN conditioning MLPs.
@@
Table 2 compares four decoding strategies across all three encoders on 10 languages. Cosine similarity is averaged over the same sentence translated into English, Chinese, German, Japanese, French, Spanish, Korean, Russian, Arabic, and Portuguese. Sequential greedy consistently achieves the highest similarity across encoders.
@@ 5 Conclusion
We presented embedding inversion via conditional masked diffusion, achieving 81.3% token accuracy across three embedding models with a 78M parameter decoder that requires no access to the target encoder. The progression from Vec2Text, which requires a compatible encoder architecture, to ALGEN, which needs alignment training, to Zero2Text, which queries the target API, to our work, which needs none of these, demonstrates that inversion attacks are becoming more accessible, not less. Production systems that cache or transmit embedding vectors under the assumption of irreversibility should treat them as sensitive data requiring protection equivalent to the original text.
Current limitations include the restriction to 32-token sequences, whereas real documents are substantially longer. Scaling to longer sequences via hierarchical diffusion or sliding window approaches is a natural next step. Incorporating language model priors through classifier-free guidance could further improve reconstruction quality, particularly for low-frequency tokens. On the defense side, encrypted vector search and inversion-resistant embeddings through adversarial training or differential privacy on the embedding space remain important open problems for security-critical deployments.